A car dealer sells 3 types of car. On the first day, he sells 3 cars of model A and 2 of model B for a sum of $260 000. On the second day, he sells 3 cars of model B and 2 cars of model C for a sum of $160 000, and on the third day, he sells 3 cars of model C and 2 cars of model A for a total of $180 000. What is the price of one model C car?
Answer by stanbon(75887)

A car dealer sells 3 types of car.
On the first day, he sells 3 cars of model A and 2 of model B for a sum of $260 000.
On the second day, he sells 3 cars of model B and 2 cars of model C for a sum of $160 000, and
on the third day, he sells 3 cars of model C and 2 cars of model A for a total of $180 000.
What is the price of one model C car?
-----------------------
Equations:
3a + 2b + 0c = 260000
0a + 3b + 2c = 160000
2a + 0b + 3c = 180000
------------------------
I used a matrix function on a TI calculator to get:
a = 60000
b = 40000
c = 20000
==============
Cheers,
Stan H.